The Comic Book Guy weighs 435 pounds and is on a diet to lose 5 pounds each week. Professor Frink weighs 123 pounds and is on a diet to gain three pounds each week. How long will it take for them to weigh the same?
Answer by robertb(5830)

435 + (n - 1)(-5) = 123 + (n-1)3;
435 - 5n + 5 = 123 + 3n - 3;
440 - 5n = 120 + 3n;
320 = 8n,
n = 40,
n - 1 = 39, the number of days after which they will weigh the same.
(The Simpsoooon's...taaantantantanantantatananananan.... =D)